OpenStudy (anonymous):

The intensity of illumination, I, produced by a light of power P at a point distant x m from the light is given by \[I =frac{ P }{ x^2 }\] Two lights, of powers P and 2P are 20 metres apart. Find to the nearest metre the distance from the less powerful light at which the intensity of illumination is least.

OpenStudy (anonymous):

\[I=\frac{ P }{ x^2 }\]

OpenStudy (anonymous):

Okay Hero.

OpenStudy (anonymous):

Thanks for that.

OpenStudy (whpalmer4):

Is there some assumption not stated about where the measurement point might be? For example, are we looking for a spot somewhere on the line between the two sources?

OpenStudy (anonymous):

Yeah I think so. I'm not particularly sure about the wording of the question. That's why I'm asking. But I think it's a spot between the two light sources.

OpenStudy (anonymous):

what happens when two lights illuminate the same spot? do we take the average of the intensity at that spot?

OpenStudy (anonymous):

do we add them together

OpenStudy (anonymous):

i would assume we would add them but this is a physics problem and there is probably something i'm missing

OpenStudy (anonymous):

if its just the summation then i would suggest minimizing the equation \[\frac{ 2P }{ (20 - x)^{2} } + \frac{ P }{ x^{2} }\]

OpenStudy (anonymous):

It's not a Physics question. It looks like one, but it's not.

OpenStudy (anonymous):

i would assume then that the intensity at any given point is the summation of the two intensities. if so then \[I _{1} = \frac{ 2P }{ (20 - x )^{2}}\] that is the brightest intensity bulb and the other would would be \[I _{2} = \frac{ P }{ x^{2} }\] where x is the distance from the bulb I assume then that the combined intensity at any point is just I1 + I2 I suggest to minimize I1 + I2 using differentiation

OpenStudy (anonymous):

So we add the two and then differentiate? Because if so, don't think that's going to work.

OpenStudy (anonymous):

yeah you'll get \[-\frac{ 2P }{ x^{3} } - \frac{ 4P }{ (x - 20)^{3} }\] then you can set this = 0

OpenStudy (anonymous):

Both aren't negative. One's got to be positive, if you move them to one side.

OpenStudy (anonymous):

But I don't think that's how you do it. I will wait for whpalmer's reply.

OpenStudy (whpalmer4):

Isn't the spot where the minimum illumination occurs going to be the spot where the illumination from each source is equal?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

the one with the less power which is the one that's P.

OpenStudy (anonymous):

oh, well if you solve for the derivative = 0 you will get x = about 8.84989 meters from the less intense bulb

OpenStudy (anonymous):

That's right!!!!!

OpenStudy (anonymous):

and this is a verified minimum because values close to that result in a higher intensity

OpenStudy (anonymous):

WTF...It's 9m

OpenStudy (anonymous):

Can you get it to exxactly 9?

OpenStudy (whpalmer4):

nearest meter

OpenStudy (anonymous):

exactly*

OpenStudy (anonymous):

oh, just round up then

OpenStudy (anonymous):

Don't think the answers rounded up. Not sure, but unlikely it will need rounding up. Because if it did, the question would tell you.

OpenStudy (whpalmer4):

question asked for "to the nearest metre"

OpenStudy (anonymous):

idk, that's what i get

OpenStudy (anonymous):

In my previous question, I had $403, but the answers had $400. So I did it again, and then found out something about the question, and redid it using something else. Got it to $400.

OpenStudy (anonymous):

Wait hold on for a minute a two there.

OpenStudy (anonymous):

Ah okay yeah you're right.

OpenStudy (anonymous):

Sorry about that guys. Thanks mate.

OpenStudy (anonymous):

Thanks both of yous. can't give you guys both medals. Lolz....

OpenStudy (whpalmer4):

Fun problem!

OpenStudy (anonymous):

@whpalmer4 @binarymimic Holy cow, I couldn't get it. Could you show me the working?

OpenStudy (anonymous):

How the hell do you factor x^6? That's where I ended when I made the derivative=0

OpenStudy (anonymous):

The derivative of P/x^2 is \[-\frac{ 2P }{ x^3 }\] and the derivative of 2P/(20 - x)^2 is \[\frac{ 4P }{ (20 - x)^3 }\]

OpenStudy (anonymous):

You forgot the minus.

OpenStudy (anonymous):

I'm stuck at when you make the whole thing equal to zero,

OpenStudy (anonymous):

minus?

OpenStudy (anonymous):

\[-\frac{ 4P }{ (20-x)^3 }\]

OpenStudy (anonymous):

oh wait nvm

OpenStudy (anonymous):

You're correct. But I'm stuck at when you make them equal to zero.

OpenStudy (anonymous):

And isn't it 4Px*

OpenStudy (anonymous):

not 4P?

OpenStudy (anonymous):

wait nvm that either lolz. I'm at a total mess here.

OpenStudy (anonymous):

we want to take their sum and equal it to 0 so \[\frac{ 4P }{ (20 - x)^{3} } -\frac{ 2P }{ x^{3} } = 0\]

OpenStudy (anonymous):

Yeah I did that, Do you combine it?

OpenStudy (anonymous):

You could set up the equality 4P * x^3 = 2P * (20 - x)^3

OpenStudy (anonymous):

Yeah but how do you factorise a cubic polynomial?

OpenStudy (anonymous):

I can't seem to factorise the cubic polynomial*

OpenStudy (anonymous):

we dont have to factor anything really first cancel the Ps and divide by 2 to get \[2x^{3} = (20 - x)^{3}\]

OpenStudy (anonymous):

see where we're going ?

OpenStudy (anonymous):

Yeah you cube root everything. So you get cube root2x which equals 20-x?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Ah okay, so you get 8.84?

OpenStudy (anonymous):

8.85*

OpenStudy (anonymous):

yes thats what i get

OpenStudy (anonymous):

How do you know it's the minimum distance? Would you have to differentiate the derivative?

OpenStudy (anonymous):

i just tested values on the original equation close to that point and found that if i used a value smaller than 8.85 then the combined intensity was larger, and if i used a value larger than 8.85, the combined intensity was still larger thats why i concluded it was a minimum

OpenStudy (anonymous):

Ah okay.

OpenStudy (anonymous):

If you want to take a second derivative then .. . >_>

OpenStudy (anonymous):

lolz, no way man. That's gonna take a trillion years.

OpenStudy (anonymous):

i used wolfram to do it, and then evaluated it at 8.849866680, and it returned a positive value. that means it is concave up at that point, so still a minimum :D

OpenStudy (anonymous):

ah okay. COuld you check this closed question how the answers got when f increases x<0?